# Band Structure and Pairing Nature of La3Ni2O7 Thin Film at Ambient Pressure

### v2

@@ Abstract
Recently, evidences of superconductivity (SC) with onset $T_{c}$ above the McMillan limit have been detected in the La3Ni2O7 ultrathin film grown on the LaSrAlO4 substrate at ambient pressure. This progress opens a new era in the field of the nickelate superconductors. Here we perform a density-functional-theory (DFT) based calculation for the band structure of this material. The obtained DFT+$U$ band structure has the feature that the bonding $d_{z^{2}}$ band crosses the Fermi level, forming the hole pocket $\gamma$, consistent with the angle-resolved-photo-emission-spectrum (ARPES). Taking the low-energy Ni-$(3d_{z^{2}},3d_{x^{2}-y^{2}})$ orbitals placed on the tetragonal lattice structure, we construct a 2D bilayer four-band tight-binding model which well captures the main features of the DFT+$U$ band structure. Then considering the multi-orbital Hubbard interaction, we adopt the random-phase approximation (RPA) approach to investigate the pairing nature. The obtained pairing symmetry is $s^{\pm}$ or $d_{xy}$ for the hole-doping level $\delta$ below or above 0.12, induced by the different Fermi surface nesting situations. For the realistic $\delta=0.21$ measured by the ARPES, our RPA calculations obtain the next-nearest-neighbor pairing $d_{xy}$-wave SC dominated by the $d_{z^{2}}$ orbital, consistent with the experimental observation that the $T_{c}$ enhances with the shrinking of the in-plane lattice constants and is insensitive to the c-axis one. This pairing state is induced by the nesting between the different patches within the $\gamma$ pocket. Our results appeal for experimental verifications.
@@ Introduction:
The discovery of high-temperature superconductivity (SC) in the Ruddlesden-Popper (RP) phase multilayer nickelate superconductors La3Ni2O7 [1] and La4Ni3O10 [2, 3, 4, 5] under high pressure (HP) has aroused a surge in the exploration of the pairing mechanism and physical properties of the nickelates family both experimentally [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] and theoretically [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 25, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 26, 27, 62, 63, 64, 28, 29, 65, 30, 31, 32, 66, 33, 67, 68, 69, 70, 71, 34, 35, 72, 73, 36, 74, 75, 37, 38, 39, 40, 41, 42, 43, 76, 44, 77, 45, 78, 79, 80, 81, 82, 46, 83, 84]. As a new platform to study the high-temperature SC, the nickelates family is different from the cuprates and the iron-based superconductors family in their multilayer structure, which is believed to play a crucial role in the pairing mechanism [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Previously, the SC in these RP phase nickelates only emerges under HP, while most experiments can only be conducted at ambient pressure (AP) due to technical difficulties. This strongly hinders the experimental investigation of the pairing mechanism of these superconductors. Very recently, this research field witnessed a breakthrough, i.e. the detection of SC with onset $T_{c}$ above the McMillan limit ($\approx$40 K) in the La3Ni2O7 [85], La2.85Pr0.15Ni2O7 [86] and La2PrNi2O7 [87] thin films grown on the LaSrAlO4 (LSAO) substrate at AP by two different teams.
@@
While the zero resistivity and Meissner effect detected by both teams identify the presence of real SC, the Brezinskii-Kosterlitz-Thouless transition character and the anisotropic critical magnetic fields reveal the 2D characteristic of the SC in these bilayer nickelate films [85, 86, 87]. Experimental tools such as the scanning transmission electron microscopy (STEM) have revealed that these film materials host the same tetragonal crystal structure as that of the bulk materials under HP [88, 89], in which there is no tilting of the NiO6 octahedra. Through tuning the strength of the strain, it is found that the $T_{c}$ enhances with the shrinking of in-plane lattice constant and is insensitive to the c-axis one [85, 88]. In the aspect of electronic structure, the angle-resolved-photo-emission-spectrum (ARPES) identifies a band structure qualitatively consistent with the density-functional-theory (DFT) $+U$ calculations [89, 90]. Notably, the bonding $d_{z^{2}}$ band top crosses the Fermi level, forming a hole pocket $\gamma$ centering around the $M(\pi,\pi)$-point similar with the case in the bulk La3Ni2O7 under HP. Significantly, from the area enclosed by the Fermi surfaces (FSs) measured by ARPES, it is found that hole-doping has been introduced to the film, probably through Sr diffusion from the substrate [90]. Currently, the pairing nature of this material remain unknown.
@@
In this paper, we construct a Ni-$(3d_{z^{2}},3d_{x^{2}-y^{2}})$ orbital bilayer tight-binding (TB) model to investigate the pairing nature of the La3Ni2O7 ultrathin film. Our DFT+U band structure shows that the bonding-$d_{z^{2}}$ band crosses the Fermi level, forming the $\gamma$-pocket, consistent with the ARPES observation. Using the two low-energy Ni-$3d$-$e_{g}$ orbitals placed in the tetragonal structure, we construct a 2D four-band TB model, which well fits the DFT band structure. After considering the multi-orbital Hubbard interactions, we engage the random-phase-approximation (RPA) approach to investigate the pairing symmetry. Our RPA results suggest that the pairing symmetry is $s^{\pm}$ or $d_{xy}$ for the hole-doping level $\delta$ below or above 0.12. For the realistic $\delta=0.21$ measured by ARPES, the $d_{xy}$-wave pairing is the leading pairing symmetry, which is induced by the nesting within the $\gamma$-pocket. The real-space pairing pattern is dominated by the next-nearest-neighbor (NNN) pairing between the $d_{z^{2}}$ orbitals. Our results appeal for experimental verifications.
[…]
DFT Band Structure:
@@
As seen in Fig. 1(a), the ultrathin film of La3Ni2O7 crystallizes in a tetragonal phase at AP, in which there is no tilting of the NiO6 octahedron. We consider a single-bilayer. First-principles DFT calculations were implemented utilizing the Vienna ab initio simulation package (VASP) and employing the projector augmented wave (PAW) as the pseudopotentials [91, 92, 93]. The electronic correlations were considered by the generalized gradient approximation (GGA) and the Perdew-Burke-Ernzerhof (PBE) exchange potential [94]. The plane-wave cutoff energy was set as 550 eV and a $k$-point grid 12 × 12 × 1 was adopted for the single-bilayer structure of La3Ni2O7 with the $p4/mm$ symmetry reported in Ref. [89], where $a=b=3.7544\ \mathrm{\AA}$ and the spacing between the two Ni-O layers is $4.28\ \mathrm{\AA}$. We performed DFT+$U$ band calculations with a reasonable onsite Coulomb interaction $U$ = 3.5 eV [1, 22].
@@
For convenience in the subsequent studies, we construct a Wannier model with the Ni-$(d_{x^{2}-y^{2}},d_{z^{2}})$ orbitals based on the maximally-localized Wannier function method implemented in the WANNIER90 code [95]. The hopping parameters in the TB model are obtained by Wannier downfolding the band structure with Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals. As shown in Fig. S1 in the Appendix A, the Wannier bands align remarkably well with the DFT bands near the Fermi level.
TB Model and Microscopic Hamiltonian:
Considering the hoppings up to a few neighborings, we obtain the following TB model,
Here $i/j$ labels site, $\mu/\nu$ labels layer, $\alpha/\beta$ labels orbital, $\sigma$ labels spin. The $t_{i\mu\alpha,j\nu\beta}$ represent for the hopping integrals, whose definition are illustrated in Fig. 2 and the corresponding values are listed in Table SI in Appendix B. The analytical formula of the TB Hamiltonian is also provided in Appendix B. In comparison with the TB parameters for bulk La3Ni2O7 at AP [24], the intralayer nearest-neighbor hopping of the $d_{x^{2}-y^{2}}$ electrons $t_{1}^{x}$ becomes stronger due to the shrinked in-plane lattice constants, while the interlayer hopping of the $d_{z^{2}}$ electrons $t_{\perp}^{z}$ gets weaker due to the elongated c-axis one.
Figure 3: (color online) Band structure and FSs of La3Ni2O7 film at AP obtained from the four-band TB-model Eq. (1). (a) The band structure of model Eq. (1) (red lines) along the high symmetry lines, compared with the DFT band structure (black lines). The blue and purple dashed lines represent the Fermi levels for the hole-doping level $\delta=0.1$ and $0.2$, respectively. (b-c) FSs in the BZ with hole-doping levels $\delta=0.1$ (b) and $\delta=0.2$ (c). In (b-c), The FSs marked as $\alpha$, $\beta$, $\gamma$ and $\eta$. The color indicates the orbital weight of $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$. The FS-nesting vector is marked by $\bm{Q_{1}}/\bm{Q_{2}}$ in (b/c).
The band structure of our four-band TB model Eq. (1) is shown in Fig. 3(a), in comparison with the DFT one. Clearly, the former has captured the main features of the latter. As the ARPES result suggests that the system is hole-doped [90], we show the FSs of two typical hole-doping levels $\delta=0.1$ and $0.2$ in Fig. 3 (b) and (c), which correspond to the average particle number per site $n=1.5-\delta=1.4$ and $1.3$, respectively. In both FSs, there exist electron pockets $\alpha$ and $\eta$ centering around the $\Gamma$ point and hole pockets $\beta$ and $\gamma$ centering around the M point. The $\alpha$ and $\beta$ pockets show mixing of orbital contents while the $\gamma$ and $\eta$ pockets are dominated by the $d_{z^{2}}$ orbital. Although the FS topologies of the two dopings are the same, their FS-nesting vectors are distinct due to their different pocket sizes. For $\delta=0.1$ the dominant nesting is between the $\eta$ and $\gamma$ pockets with nesting vector $\bm{Q}_{1}\approx(0.67\pi,0.67\pi)$; while for $\delta=0.2$ the dominant nesting is between the different parts of the $\gamma$ pocket with nesting vector $\bm{Q}_{2}\approx(0.93\pi,0.2\pi)$.
$\left.+\left(c^{{\dagger}}_{i\mu{}z\uparrow}c^{{\dagger}}_{i\mu{}z\downarrow}c_{i\mu{}x\downarrow}c_{i\mu{}x\uparrow}+\mathrm{h.c.}\right)\right].$
Here the $U$ and $V$ terms are the intraorbital and interorbital repulsions, and the $J_{H}$ term includes the Hund’s coupling and pair hopping. We use the relation $U=V+2J_{H}$ and set $J_{H}=U/6$ in our study.
Figure 4: (color online) (a) The leading pairing symmetry as function of the hole-doping level $\delta=1.5-n$. (b-c) The RPA results for $\delta=0.1$. (b) The distribution of the spin susceptibility $\chi^{(s)}(\bm{k})$ over the BZ for $U<U_{c}$. The maximal value of $\chi^{(s)}(\bm{k})$ just locates at the FS-nesting vector $\bm{Q_{1}}\approx(0.67\pi,0.67\pi)$. (c) The largest pairing eigenvalue $\lambda$ as function of $U$ for different pairing symmetries. (d-e) The RPA results for $\delta=0.2$. (d) The distribution of the spin susceptibility $\chi^{(s)}(\bm{k})$ over the BZ for $U<U_{c}$. The maximal value of $\chi^{(s)}(\bm{k})$ also locates at the FS-nesting vector $\bm{Q_{2}}\approx(0.93\pi,0.2\pi)$. (e) The largest pairing eigenvalue $\lambda$ as function of $U$ for different pairing symmetries.
Doping-Dependent Pairing Symmetries:
The Hamiltonian Eq. (2) is solved by standard multi-orbital RPA approach  [96, 97, 98, 99, 100, 101, 102]. In the RPA framework, the SC is driven by the spin fluctuation whose propagator is given by the spin susceptibility renormalized up to the RPA level. When $U$ is below the critical interaction strength $U_{c}$, the $T_{c}$ of the SC mediated by the spin fluctuation is determined by the pairing eigenvalue $\lambda$ through $T_{c}\sim e^{-1/\lambda}$, and the pairing symmetry is determined by the corresponding pairing eigenvector.
For the case of $\delta\leq 0.12$, we present the results in Fig. 4 (b) and (c) for a typical doping level $\delta=0.1$. For this doping, the distribution of the spin susceptibility $\chi^{(s)}(\bm{q})$ over the BZ for $U<U_{c}$ is shown in Fig. 4 (b). The distribution is just maximized at the FS-nesting vector for this doping, i.e. $\bm{Q}_{1}\approx(0.67\pi,0.67\pi)$, which is between the $\eta$ and $\gamma$ pockets. Fig. 4 (c) shows the $U$-dependent $\lambda$ for the various leading pairing symmetries. The $\lambda$ for these pairing symmetries enhance promptly with the enhancement of $U$ for $U<U_{c}$. The leading and subleading pairing symmetries are the $s$-wave and the $d_{xy}$-wave for all the $U$ parameters, respectively. The pairing gap function of the leading $s$-wave pairings changes sign between the nested patches on the $\eta$- and $\gamma$- pockets, leading to an $s^{\pm}$-wave pairing, see Fig. S2 in Appendix C.
For the case of $\delta>0.12$, we present the results in Fig. 4 (d) and (e) for a typical doping level $\delta=0.2$. For this doping, the distribution of the spin susceptibility $\chi^{(s)}(\bm{q})$ over the BZ for $U<U_{c}$ is shown in Fig. 4 (d). The distribution is also maximized at the corresponding FS-nesting vector, $\bm{Q}_{2}\approx(0.93\pi,0.2\pi)$, which is between the different parts of the $\gamma$ pocket. Fig. 4 (e) shows the $U$-dependence of $\lambda$ for different pairing symmetries, which suggests that the leading and subleading pairing symmetries are the $d_{xy}$-wave and $s$-wave, respectively.
The $d_{xy}$-Wave Pairing:
Since the ARPES measurement suggests the hole-doping level to be about $0.21$, we focus on the typical case of $\delta=0.2$ hereby. The distribution of the gap function of the leading pairing state on the FSs for $\delta=0.2$ is shown in Fig. 5(a). In Fig. 5 (a), the gap function changes sign with every 90o rotation. It also changes sign upon mirror reflection about the $xz$ or the $yz$ plane, giving rise to gap nodes on the x- or y-axes. Such a gap function belongs to the $d_{xy}$-wave pairing symmetry. In this pairing state, the strongest pairing amplitude locates on the patches connected by the nesting vector $\bm{Q}_{2}$ within the $\gamma$-pocket, with the gap signs on these nested patches opposite to each other, as shown in Fig. 5 (a). Such a gap-sign structure maximizes the energy gain by the superconducting condensation. In contrast, the $s$-wave pairing symmetry is energetically disfavored here because it can hardly require that the patches connected by the nesting vector $\bm{Q}_{2}$ carry opposite gap signs, see Fig. S2 in Appendix C. Note that, different from the cases in the bulk La3Ni2O7 under HP [1] and the iron-based superconductors, the FS-nesting here is not between different pockets but between the different patches within one pocket, similar with the case in the cuprates. Such FS-nesting can also favor SC when the nested patches are straight enough, see Fig. 5 (a).
Figure 5: (color online) (a) Distributions of the leading $d_{xy}$-wave pairing gap functions on the FSs for $\delta=0.2$ and $U=0.75$ eV. $\bm{Q_{2}}$ marks the FS-nesting. (b) Schematic of the pairing in the real space for $d_{xy}$-wave pairing.
For convenience in presenting the real-space pairing configuration, we illustrate the few dominant pairing amplitudes on the few neighboring bonds in Fig. 5(b). Our RPA results suggest that all interorbital pairing components are negligible and thus all the shown pairing components are intra-orbital ones. For intraorbital pairings, constrained by the $D_{4}$ point group symmetry, the interlayer pairing on the vertical bonds and the intralayer pairing on the nearest-neighbor bonds are forbidden by the $d_{xy}$ pairing symmetry. Our RPA results yield that the strongest pairing component is the intralayer next-nearest-neigbor (NNN) pairing for the $d_{z^{2}}$-orbital, i.e. the $\Delta_{z}$ shown in Fig. 5(b). The reason for this lies in two aspects. Firstly, as shown in Fig. 5 (a), the dominant pairing amplitude is located on the $\gamma$ pocket, which is dominated by the $d_{z^{2}}$ orbital. Secondly, the nearest pairing component allowed by the $d_{xy}$-wave pairing symmetry is the NNN one. The next strongest pairing component is the interlayer NNN pairing for the $d_{z^{2}}$-orbital, i.e. the $\Delta_{z}^{\prime}$ shown in Fig. 5(b), with $\Delta_{z}^{\prime}/\Delta_{z}=0.9$. The fact that the intralayer and interlayer NNN pairing amplitudes are nearly equal originates from that the strongest pairing amplitude shown in Fig. 5 (a) is located on the bonding $d_{z^{2}}$ band in which the orbital weights from both layers are equally mixed. In addition, there are weak intralayer and interlayer NNN pairing components for the $d_{x^{2}-y^{2}}$ orbital, i.e. $\Delta_{x}=0.08\Delta_{z}$ and $\Delta_{x}^{\prime}=0.01\Delta_{z}$. All the other pairing components are negligibly weak.
Discussion and Conclusions:
The band feature shown in Fig. 3(a) that the bonding-$d_{z^{2}}$ band crosses the Fermi level is consistent with the ARPES observation. The NNN-pairing $d_{xy}$-wave SC obtained here is consistent with the fact that the $T_{c}$ enhances with the shrinking of the in-plane lattice constants and is insensitive to the c-axis one [85]. This pairing symmetry can be verified in the ARPES experiment by the presence of gap nodes on the x- and y- axes. In addition, it can be identified by phase sensitive experiments.
Acknowledgements
Appendix A Appendix A. Wannier Band Structure
We construct a Wannier model with the Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals based on the maximally-localized Wannier function method, whose band structure is shown in Fig. S1.
Figure S1: (color online) Comparison of the density-functional-theory (black solid line) and Wannier band (red solid line) structures along a high symmetry path in the Brillouin zone.
Appendix B Appendix B. Tight-Binding Hamiltonian
The values of the hopping parameters defined in the main text are listed in Table SI.
@@ Acknowledgements
Table SI: The hopping integrals of the $(d_{z^{2}},d_{x^{2}-y^{2}})$-orbital bilayer TB model for the La3Ni2O7 film at AP. In the superscript and subscript, $x(z)$ represents the $d_{x^{2}-y^{2}}(d_{z^{2}})$ orbit, $\perp$ represents interlayer hopping, and $1,2,3$ represents the nearest-neighbor, next-nearest-neighbor and third-nearest-neighbor hopping, respectively. $\epsilon$ is on-site energy. The unit of all parameters is eV.
@@
The tight-binding (TB) Hamiltonian can be transformed in the $\bm{k}$-space since the system has translation symmetry.
$H_{\mathrm{TB}}=$ $\sum_{ij\mu\nu\alpha\beta\sigma}t_{i\mu\alpha,j\nu\beta}c^{{\dagger}}_{i\mu\alpha\sigma}c_{j\nu\beta\sigma}$ (S1)
$=$ $\sum_{\bm{k}\mu\nu\alpha\beta\sigma}\left(H_{\bm{k}}\right)_{\mu\alpha,\nu\beta}c^{{\dagger}}_{\bm{k}\mu\alpha\sigma}c_{\bm{k}\nu\beta\sigma}$
Here $i/j$ labels site, $\mu/\nu$ labels layer, $\alpha/\beta$ labels orbital ($d_{z^{2}}$ ($z$) and $d_{x^{2}-y^{2}}$ ($x$)), $\sigma$ labels the spin, $t_{i\mu\alpha,j\nu\beta}$ represents for the corresponding hopping integrals and $H_{\bm{k}}$ is the $\bm{k}$-space TB Hamiltonian matrix. The effective orbital indices is the combined layer ($t$/$b$) and physical orbital indices ($x$/$z$). In the basis $(tz,tx,bz,bx)$, $H_{\bm{k}}$ takes the form
$H_{\bm{k}}=\left[\begin{array}[]{cc}H_{1}&H_{2}\\ H_{2}^{{\dagger}}&H_{1}\end{array}\right]$ (S2)

### v3

@@ Abstract
Recently, evidences of superconductivity (SC) with onset $T_{c}$ above the McMillan limit have been detected in the La3Ni2O7 ultrathin film grown on the LaSrAlO4 substrate at ambient pressure. This progress opens a new era in the field of the nickelate superconductors. Here we perform a density-functional-theory (DFT) based calculation for the band structure of this material. The obtained DFT+$U$ band structure has the feature that the bonding $d_{z^{2}}$ band crosses the Fermi level, forming the hole pocket $\gamma$, consistent with the angle-resolved photoemission spectrum (ARPES). Taking the low-energy Ni-$(3d_{z^{2}},3d_{x^{2}-y^{2}})$ orbitals placed on the tetragonal lattice structure, we construct a 2D bilayer four-band tight-binding model which well captures the main features of the DFT+$U$ band structure. Then considering the multi-orbital Hubbard interaction, we adopt the random-phase approximation (RPA) approach to investigate the pairing nature. The obtained pairing symmetry is $s^{\pm}$ or $d_{xy}$ for the hole-doping level $\delta$ below or above 0.12, induced by the different Fermi surface nesting situations. For the realistic $\delta=0.21$ measured by the ARPES, our RPA calculations obtain the next-nearest-neighbor pairing $d_{xy}$-wave SC dominated by the $d_{z^{2}}$ orbital, consistent with the experimental observation that the $T_{c}$ enhances with the shrinking of the in-plane lattice constants. This pairing state is induced by the nesting between the different patches within the $\gamma$ pocket. Our results appeal for experimental verifications.
@@ Introduction:
The discovery of high-temperature superconductivity (SC) in the Ruddlesden-Popper (RP) phase multilayer nickelate superconductors La3Ni2O7 [1] and La4Ni3O10 [2, 3, 4, 5] under high pressure (HP) has aroused a surge in the exploration of the pairing mechanism and physical properties of the nickelates family both experimentally [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] and theoretically [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 25, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 26, 27, 62, 63, 64, 28, 29, 65, 30, 31, 32, 66, 33, 67, 68, 69, 70, 71, 34, 35, 72, 73, 36, 74, 75, 37, 38, 39, 40, 41, 42, 43, 76, 44, 77, 45, 78, 79, 80, 81, 82, 46, 83, 84, 85]. As a new platform to study the high-temperature SC, the nickelates family is different from the cuprates and the iron-based superconductors family in their multilayer structure, which is believed to play a crucial role in the pairing mechanism [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Previously, the SC in these RP phase nickelates only emerges under HP, while most experiments can only be conducted at ambient pressure (AP) due to technical difficulties. This strongly hinders the experimental investigation of the pairing mechanism of these superconductors. Very recently, this research field witnessed a breakthrough, i.e. the detection of SC with onset $T_{c}$ above the McMillan limit ($\approx$40 K) in the La3Ni2O7 [86], La2.85Pr0.15Ni2O7 [87] and La2PrNi2O7 [88] ultrathin films grown on the LaSrAlO4 (LSAO) substrate at AP by two different teams.
@@
While the zero resistivity and Meissner effect detected by both teams identify the presence of real SC, the Brezinskii-Kosterlitz-Thouless transition character and the anisotropic critical magnetic fields reveal the 2D characteristic of the SC in these bilayer nickelate films [86, 87, 88]. Experimental tools such as the scanning transmission electron microscopy (STEM) have revealed that these film materials host the same tetragonal crystal structure as that of the bulk materials under HP [89, 90], in which there is no tilting of the NiO6 octahedra. Through tuning the strength of the strain, it is found that the $T_{c}$ enhances with the shrinking of in-plane lattice constant and is insensitive to the c-axis one [86, 89]. In the aspect of electronic structure, the angle-resolved photoemission spectrum (ARPES) identifies a band structure qualitatively consistent with the density-functional-theory (DFT) +U+U calculations [90, 91]. Notably, the bonding $d_{z^{2}}$ band top crosses the Fermi level, forming a hole pocket $\gamma$ centering around the $M(\pi,\pi)$-point similar with the case in the bulk La3Ni2O7 under HP. Significantly, from the area enclosed by the Fermi surfaces (FSs) measured by ARPES, it is found that hole-doping has been introduced to the film, probably through Sr diffusion from the substrate [91]. Currently, the pairing nature of this material remains unknown.
@@
In this paper, we construct a Ni-$(3d_{z^{2}},3d_{x^{2}-y^{2}})$ orbital bilayer tight-binding (TB) model to investigate the pairing nature of the La3Ni2O7 ultrathin film. Our DFT+U+U band structure shows that the bonding-$d_{z^{2}}$ band crosses the Fermi level, forming the $\gamma$-pocket, consistent with the ARPES observation. Using the two low-energy Ni-$3d$-$\mathrm{e_{g}}$ orbitals placed in the tetragonal structure, we construct a 2D four-band TB model, which well fits the DFT band structure. After considering the multi-orbital Hubbard interactions, we engage the random-phase-approximation (RPA) approach to investigate the pairing symmetry. Our RPA results suggest that the pairing symmetry is $s^{\pm}$ or $d_{xy}$ for the hole-doping level $\delta$ below or above 0.12. For the realistic $\delta=0.21$ measured by ARPES, the $d_{xy}$-wave pairing is the leading pairing symmetry, which is induced by the nesting within the $\gamma$-pocket. The real-space pairing pattern is dominated by the next-nearest-neighbor (NNN) pairing between the $d_{z^{2}}$ orbitals. Our results appeal for experimental verifications.
Figure 1: (color online) The DFT electronic structures. (a) Side- (upper) and top- (lower) views of the crystal structure of the La3Ni2O7 single bilayer. (b) 2D BZ with high-symmetric points and lines marked. (c) DFT+$U$ band structure of the La3Ni2O7 single-bilayer with $U$ = 3.5 eV. The blue and red colors represent for the weights of Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals respectively. (d) The DOS contributed by La, O and Ni-3dd orbitals.
DFT Band Structure and TB Model:
@@
As seen in Fig. 1(a), the ultrathin film of La3Ni2O7 crystallizes in a tetragonal phase at AP, in which there is no tilting of the NiO6 octahedron. We consider a single-bilayer. First-principles DFT calculations were implemented utilizing the Vienna ab initio simulation package (VASP) and employing the projector augmented wave (PAW) as the pseudopotentials [92, 93, 94]. The electronic correlations were considered by the generalized gradient approximation (GGA) and the Perdew-Burke-Ernzerhof (PBE) exchange potential [95]. The plane-wave cutoff energy was set as 550 eV and a $k$-point grid 12 × 12 × 1 was adopted for the single-bilayer structure of La3Ni2O7 with the $p4/mm$ symmetry reported in Ref. [90], where a=b=3.7544Åa=b=3.7544\ $\mathrm{\AA }$ and the spacing between the two NiO2 layers is 4.28Å4.28\ $\mathrm{\AA }$. We performed DFT+$U$ band calculations with a reasonable onsite Coulomb interaction $U$ = 3.5 eV [1, 22] to correct the band structure in strongly correlated electron systems.
@@
For convenience in the subsequent studies, we construct a Wannier model with the Ni-$(d_{x^{2}-y^{2}},d_{z^{2}})$ orbitals based on the maximally-localized Wannier function method implemented in the WANNIER90 code [97]. The hopping parameters in the TB model are obtained by Wannier downfolding the band structure with Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{z^{2}}$ orbitals. As shown in Fig. 3(a), the Wannier bands align remarkably well with the DFT bands near the Fermi level.
@@
Table 1: The hopping integrals of the $(d_{z^{2}},d_{x^{2}-y^{2}})$-orbital bilayer TB model for the La3Ni2O7 film at AP. In the superscript and subscript, $x(z)$ represents the $d_{x^{2}-y^{2}}(d_{z^{2}})$ orbit, $\perp$ represents interlayer hopping, and $1,2,3$ represents the nearest-neighbor, next-nearest-neighbor and third-nearest-neighbor hopping, respectively. $\epsilon$ is on-site energy. The unit of all parameters is eV.
@@
TB Model and Microscopic Hamiltonian:
Considering the hoppings up to a few neighborings, the TB model can be obtained and transformed into the $\bm{k}$-space,
$=$ $\sum_{\bm{k}\mu\nu\alpha\beta\sigma}\left(H_{\bm{k}}\right)_{\mu\alpha,\nu\beta}c^{{\dagger}}_{\bm{k}\mu\alpha\sigma}c_{\bm{k}\nu\beta\sigma}.$
Here $i/j$ labels site, $\mu/\nu$ labels layer, $\alpha/\beta$ labels orbital ($d_{z^{2}}$ ($z$) and $d_{x^{2}-y^{2}}$ ($x$)), $\sigma$ labels the spin, the $t_{i\mu\alpha,j\nu\beta}$ represent for the hopping integrals, whose definitions are illustrated in Fig. 2 and the corresponding values are listed in Tab. 1. $H_{\bm{k}}$ is the $\bm{k}$-space TB Hamiltonian matrix. The effective orbital indices is the combined layer (tt/$b$) and physical orbital indices (xx/$z$). In the basis $(tz,tx,bz,bx)$, the $k$-space TB Hamiltonian $H_{\bm{k}}$ takes the form
$H_{\bm{k}}=\left[\begin{array}[]{cc}H_{1}&H_{2}\\ H_{2}^{{\dagger}}&H_{1}\end{array}\right].$ (2)